# A Structural Characterization of the Hit Image in the Motivic Steenrod Algebra

### v3

@@ 3.2. A linear algebra lemma
We give a simple linear algebra fact that will be used repeatedly.
@@
3.3. The $M_{1}$–part of the $Q_{0}$–image
We now translate Kameko’s description of the image of $Q_{0}$ into a parity statement. Let $N_{n}=N_{n}(\mathbb{C})$ and $A^{\sharp}=A^{*,*}/(\tau)$ as before.
@@ 3.3. The $M_{1}$–part of the $Q_{0}$–image
    The quotient $(N_{n}/H_{n})_{d,*}$ decomposes as
    $(N_{n}/H_{n})_{d,*}\cong\bigoplus_{a+2b=d}\Lambda^{a}_{n}\otimes(Y_{n}/G_{n})^{2b},$
    and the summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ has basis given by the images of the monomials in $M_{0}\cup M_{1}$.
@@
    For each monomial $z\in N_{n}$ of degree d1d-1 with $\omega_{0}(z)=k+1$, the image $\pi(Q_{0}(z))\in\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ is a linear combination of $\pi(z^{\prime})$ with $z^{\prime}\in M_{0}$ and elements of the form $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$ with $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$.
@@
    There exist monotone permutations $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ (explicitly, the ones with $\sigma_{1}(1)=1,\sigma_{1}(2)=3,\dots,\sigma_{1}(k)=k+1$ and $\sigma_{2}(1)=2,\dots,\sigma_{2}(k)=k+1$) and a monomial $z\in N_{n}^{d-1,*}$ with $(\omega_{0}(z),\omega_{1}(z))=(k+1,n-2)$ and $\alpha_{i}(z)=k$ for all $i$, such that the $M_{1}$–part of $\pi(Q_{0}(z))$ is precisely $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$.
The first two items are exactly Propositions 4.1 and 5.1 in [3]; the third is a reformulation of the second case in [3, Prop. 5.3] in our notation.
We now apply Lemma 3.2.
Proposition 3.4 (Parity of the $Q_{0}$–image in $V$).
With notation as above, identify $V$ with $W=\mathbb{F}_{2}^{N}$ by mapping the basis $\pi(\sigma_{i}(z_{k}))$ to $e_{i}$. Then the $M_{1}$–part of the image of
$Q_{0}:N_{n}^{d-1,*}\to N_{n}^{d,*}$
is exactly the even-parity hyperplane
$E=\{w\in W\mid w_{1}+\cdots+w_{N}=0\}\subset W.$
Equivalently,
$\varepsilon\bigl(\pi(Q_{0}(N_{n}^{d-1,*}))\cap V\bigr)=0.$
@@
By Proposition 3.3(2), the $M_{1}$–component of $\pi(Q_{0}(z))$ is always a linear combination of pairwise sums $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$, so the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ is contained in the subspace $E\subset W$ spanned by all $e_{i}+e_{j}$.
Conversely, by Proposition 3.3(3) there exist $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ and a monomial $z_{0}\in N_{n}^{d-1,*}$ such that the $M_{1}$–part of $\pi(Q_{0}(z_{0}))$ is precisely $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$. Let $i_{0},j_{0}$ be the indices corresponding to $\sigma_{1},\sigma_{2}$, so that this $M_{1}$–part is the vector $e_{i_{0}}+e_{j_{0}}$ in $W$.
Now let $S_{n}$ act on $N_{n}$ by permuting the variables $x_{1},\dots,x_{n},y_{1},\dots,y_{n}$. This action preserves the subspaces $H_{n}$ and $G_{n}$, and it commutes with $Q_{0}$ and with the projection $\pi$, so for any permutation $\tau\in S_{n}$ we have
$\pi(Q_{0}(\tau\cdot z_{0}))=\tau\cdot\pi(Q_{0}(z_{0})).$
In particular, the $M_{1}$–part of $\pi(Q_{0}(\tau\cdot z_{0}))$ is $\pi(\tau\sigma_{1}(z_{k})+\tau\sigma_{2}(z_{k}))$, which corresponds to the vector $e_{i}+e_{j}$ where $i,j$ index the basis elements $\tau\sigma_{1}(z_{k})$ and $\tau\sigma_{2}(z_{k})$ of $V$.
As $\tau$ varies, the pairs of kk–subsets $\tau\sigma_{1}(\{1,\dots,k\})$ and $\tau\sigma_{2}(\{1,\dots,k\})$ run through all pairs of kk–subsets differing in exactly one element; that is, through all edges of the Johnson graph $J(n,k)$ on the vertex set $\mathrm{Mono}(k)$. It is well-known that the Johnson graph $J(n,k)$ is connected for $1\leq k\leq n-1,$ between any two basis elements $e_{i}$ and $e_{j}$ there is a path $i=v_{0},v_{1},\dots,v_{t}=j$ in $J(n,k)$, and in characteristic $2$ we have
$e_{i}+e_{j}=(e_{v_{0}}+e_{v_{1}})+(e_{v_{1}}+e_{v_{2}})+\cdots+(e_{v_{t-1}}+e_{v_{t}}).$
Each summand $e_{v_{\ell-1}}+e_{v_{\ell}}$ is realized as the $M_{1}$–part of some $\pi(Q_{0}(z))$ as above, so $e_{i}+e_{j}$ lies in the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$. Thus all vectors $e_{i}+e_{j}$ are contained in that image.
We conclude that the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ contains the subspace $E$ spanned by all $e_{i}+e_{j}$. Combined with the previous inclusion, this shows that it is exactly $E$.
By Lemma 3.2, $E$ is the even-parity hyperplane in $W$. In particular, $\varepsilon$ vanishes on $\pi(Q_{0}(N_{n}^{d-1,*}))\cap V$. ∎
3.4. The $A^{>0,*}$–image and the odd-parity gap
@@
We next extend Proposition 3.4 from the image of $Q_{0}$ to the image of the full motivic Steenrod algebra in the summand $V$.
Lemma 3.5 (Reduced powers do not contribute to $M_{1}$).
In the summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ of $(N_{n}/H_{n})_{d,*}$, the $M_{1}$–component of the image of the reduced powers $\mathcal{P}^{a}$ ($a\geq 1$) is zero. Equivalently, modulo $H_{n}$ and modulo $\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$, the only contribution to $V$ from $A^{\sharp}_{+}(N_{n})$ comes from $Q_{0}$.
@@
By construction, $G_{n}\subset Y_{n}$ is spanned by $P^{+}(Y_{n})$ and monomials of smaller weight, where $P$ is the subalgebra of $A^{\sharp}$ generated by the reduced powers $\mathcal{P}^{a}$ [3, Sec. 4]. In particular, $P^{+}(Y_{n})\subset G_{n}$ by definition.
For $a\geq 1$, the operation $\mathcal{P}^{a}$ has bidegree $(2a,a)$ and acts trivially on $\Lambda_{n}$, so for any monomial $x\otimes y\in\Lambda^{k}_{n}\otimes Y_{n}^{2(d_{1}-a)}\subset N_{n}^{d-2a,*}$ we have
$\mathcal{P}^{a}(x\otimes y)=x\otimes\mathcal{P}^{a}(y)\in\Lambda^{k}_{n}\otimes\bigl(P^{+}(Y_{n})\cap Y_{n}^{2d_{1}}\bigr)\subset\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}.$
Here we used that $2(d_{1}-a)+2a=2d_{1}$, so the image lands in topological degree $d$ and polynomial degree $2d_{1}$.
The summand $\Lambda^{k}_{n}\otimes(Y_{n}/G_{n})^{2d_{1}}$ of $(N_{n}/H_{n})_{d,*}$ is the image of $\Lambda^{k}_{n}\otimes Y_{n}^{2d_{1}}$ under the quotient map $\pi:Y_{n}\to Y_{n}/G_{n}$. Since $\mathcal{P}^{a}(x\otimes y)\in\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$ for all $x\otimes y$, its image under $\pi$ is zero. Thus the $M_{1}$–component of the image of every $\mathcal{P}^{a}$ ($a\geq 1$) in this summand vanishes. Equivalently, modulo $H_{n}$ and $\Lambda^{k}_{n}\otimes G_{n}^{2d_{1}}$, the only nontrivial image in $V$ of $A^{\sharp}_{+}(N_{n})$ comes from $Q_{0}$. ∎
Definition 3.6 (The $M_{1}$–projection to $V$).
Let $q:N_{n}\to N_{n}/H_{n}$ be the quotient map. In degree $d=k+2d_{1}$, Kameko’s decomposition [3, Prop. 4.1] gives a direct sum splitting of $(N_{n}/H_{n})_{d,*}$, hence a canonical projection
$\operatorname{pr}_{k,d_{1}}:(N_{n}/H_{n})_{d,*}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Let
$U_{0}:=\bigl\langle\pi(m)\mid m\in M_{0}\bigr\rangle\subset\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Since $\pi(M_{0})\cup\pi(M_{1})$ is a basis of $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$, we have a direct sum decomposition
$\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V,$
where $V=\langle\pi(M_{1})\rangle$ as in Section 3. Let
$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
be the projection with kernel $U_{0}$ (equivalently, $p_{M_{1}}$ kills $\pi(M_{0})$ and is the identity on $\pi(M_{1})$).
Finally, define the composite linear map
$\theta:N_{n}^{d,*}\xrightarrow{\ q\ }(N_{n}/H_{n})_{d,*}\xrightarrow{\ \operatorname{pr}_{k,d_{1}}\ }\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\xrightarrow{\ p_{M_{1}}\ }V.$
We call $\theta(u)$ the $M_{1}$–component (in $V$) of $u$.
Combining Proposition 3.4 and Lemma 3.5, we obtain the following.
Proposition 3.7 (Parity vanishes on the $A^{\sharp}_{+}$–image in the $M_{1}$–summand).
Let $u\in N_{n}^{d,*}$ be hit, i.e. $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$. Then the $M_{1}$–component $\theta(u)\in V$ has even parity:
$\varepsilon(\theta(u))=0.$
Equivalently,
$\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)\subseteq\ker(\varepsilon)\subset V.$
Proof.
By Lemma 3.5, in the summand $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ the reduced powers $\mathcal{P}^{a}$ ($a\geq 1$) contribute trivially to the $M_{1}$–part. Therefore, after applying the projection $\theta$, the only possible contribution from $A^{\sharp}_{+}(N_{n})$ comes from the Bockstein $Q_{0}$.
By Proposition 3.4, the $M_{1}$–part of $\pi(Q_{0}(N_{n}^{d-1,*}))$ is contained in the even-parity hyperplane $\ker(\varepsilon)\subset V$. Since $\theta$ is exactly the operation of passing to $(N_{n}/H_{n})_{d,*}$, projecting to the $(k,d_{1})$–summand, and then extracting the $M_{1}$–component, we obtain $\varepsilon(\theta(u))=0$ for every hit element $u$ of degree $d$. ∎
@@
Theorem 3.8 (Theorem 1.4, General odd-parity non-hit theorem).
Let $n\geq 2$ and $1\leq k<n$, and let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. Let $M_{1}$ and $V$ be as in (3.1), and let $\varepsilon:V\to\mathbb{F}_{2}$ be the parity functional from Definition 3.1. Let $\theta:N_{n}^{d,*}\to V$ be the $M_{1}$–projection from Definition 3.6. Then:
@@
    $\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
@@
    Consequently, $V/\theta(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*})$ is 11–dimensional over $\mathbb{F}_{2}$, generated by the class of any odd-parity element of $V$.
@@
    In particular, if $u\in\langle M_{1}\rangle\subset N_{n}^{d,*}$ has odd parity in its $M_{1}$–coordinates, then $u\notin A^{\sharp}_{+}(N_{n})$, so $u$ represents a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$ and hence in $QM_{n}^{d,*}(\mathbb{C})$.
@@
By Proposition 3.7, we have the inclusion
$\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)\subseteq\ker(\varepsilon).$
For the reverse inclusion, Proposition 3.4 shows that for any $i\neq j$ the vector $e_{i}+e_{j}$ (i.e. any pairwise sum $\pi(\sigma_{i}(z_{k})+\sigma_{j}(z_{k}))$) occurs as the $M_{1}$–part of $\pi(Q_{0}(z))$ for some $z\in N_{n}^{d-1,*}$, hence lies in $\theta(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*})$. These vectors span $\ker(\varepsilon)$ by Lemma 3.2. Therefore,
$\ker(\varepsilon)\subseteq\theta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr),$
which proves (1).
Since $\ker(\varepsilon)$ has codimension $1$ in $V$, statement (2) follows.
Finally, if $u\in\langle M_{1}\rangle$ has odd parity, then $\theta(u)=\pi(u)\in V$ also has odd parity (because $\theta$ restricts to the identity on $\pi(M_{1})$ by construction). If $u$ were hit, then $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$ would imply $\theta(u)\in\ker(\varepsilon)$ by (1), contradicting odd parity. Hence $u$ is not hit and defines a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$, and therefore in $QM_{n}^{d,*}(\mathbb{C})$. ∎
@@
Let $\bar{u}_{S}$ be the image of $u_{S}$ in $QN_{n}^{d,*}(\mathbb{C})$ and hence in $QM_{n}^{d,*}(\mathbb{C})$.
  1. (1)
    If $|S|$ is odd, then $\bar{u}_{S}$ is non-zero.
  2. (2)
    If $|S|$ is even, then the $M_{1}$–part of $\bar{u}_{S}$ lies in the image of $A^{\sharp}_{+}$, i.e. $\varepsilon(\bar{u}_{S})=0$.
In particular, each individual translate $\sigma(z_{k})$ represents a non-zero class, and any sum of an odd number of translates does as well.
@@
The $M_{1}$–component of $u_{S}$ is precisely $\sum_{\sigma\in S}\pi(\sigma(z_{k}))$, whose parity is $|S|$ modulo $2$. If $|S|$ is odd, this parity is $1$, so by Theorem 3.8 the class of $u_{S}$ is non-zero in the hit quotient. If $|S|$ is even, the parity is 0 and the $M_{1}$–component lies in $\ker(\varepsilon)$, which is the $M_{1}$–part of the $A^{\sharp}_{+}$–image. ∎
4. Arithmetic of the Kameko degrees
In this section we study the function $\beta(d)$ for the degrees $d=(n-1)(2^{k+1}-2)+k$ which arise in Kameko’s construction. The starting point is Proposition 2.1, which relates $\beta(d)>n$ to the inequality $\alpha(d+n)>n$.
4.1. Kameko’s original estimates
We recall Kameko’s computation from [3, Remark 2.2 and Proof of Prop. 1.5].
Proposition 4.1 (Kameko [3, Remark 2.2]).
Let $d=(n-1)(2^{k+1}-2)+k$.
  1. (1)
    If k=n1k=n-1, then $d+n=(n-1)\cdot 2^{n}+1$ and $\alpha(d+n)=\alpha(n-1)+1\leq n$.
  2. (2)
    If k=n2k=n-2, then $d+n=(n-1)\cdot 2^{n-1}$ and $\alpha(d+n)=\alpha(n-1)\leq n$.
Hence $\beta(d)\leq n$ for k=n1k=n-1 and k=n2k=n-2.
Proposition 4.2 (Kameko [3, Prop. 1.5]).
Let $d=(n-1)(2^{k+1}-2)+k$ with k=n3k=n-3. Then
$d+n=(n-1)\cdot 2^{n-2}-1=(n-2)\cdot 2^{n-2}+2^{n-3}+\cdots+2^{0}.$
If $\alpha(n-2)\geq 3$, then $\alpha(d+n)>n$, and hence $\beta(d)>n$.
Thus Kameko’s original family of motivic Peterson counterexamples arises from the case k=n3k=n-3. We now show that there are further families with k=n4k=n-4.
4.2. A new family with k=n4k=n-4
We continue to set $d=(n-1)(2^{k+1}-2)+k$ and $d_{1}=(n-1)(2^{k}-1)$, so that $d=k+2d_{1}$.
Lemma 4.3.
Let c=nkc=n-k. Then $1\leq c\leq n-1$ and
$d+n=(n-1)\cdot 2^{n-c+1}+(2-c).$
In particular, when $c=4$ (i.e. k=n4k=n-4) one has
$d+n=(n-1)\cdot 2^{n-3}-2.$
Proof.
By (2.3) and k=nck=n-c we have
$d$ $=(n-1)\bigl(2^{k+1}-2\bigr)+k$
$=(n-1)\bigl(2^{n-c+1}-2\bigr)+(n-c)$
$=(n-1)\cdot 2^{n-c+1}-2(n-1)+n-c.$
Adding $n$ gives
$d+n=(n-1)\cdot 2^{n-c+1}-2(n-1)+2n-c=(n-1)\cdot 2^{n-c+1}+(2-c),$
as claimed. For $c=4$ this simplifies to
$d+n=(n-1)\cdot 2^{n-3}-2.$
We now specialize to a particular infinite sequence of $n$.
Theorem 4.4 (Theorem 1.6, New numerical family with $\beta(d)>n$).
Let $r\geq 5$ and set $n=2^{r}+1$ and k=n4k=n-4. Let $d=(n-1)(2^{k+1}-2)+k$ as above. Then $\beta(d)>n$.
Proof.
We use Proposition 2.1. Set c=nkc=n-k. Since k=n4k=n-4, we have $c=4$, so by Lemma 4.3 we obtain
$d+n=(n-1)\cdot 2^{n-3}-2.$
Now assume $n=2^{r}+1$ with $r\geq 1$, so $n-1=2^{r}$. Then
$d+n=2^{r}\cdot 2^{n-3}-2=2^{r+n-3}-2.$
Let L=r+n3L=r+n-3. The binary expansion of $2^{L}-1$ is a string of $L$ ones, so $2^{L}-2=(2^{L}-1)-1$ has binary expansion $11\cdots 10$ with L1L-1 ones. Thus
$\alpha(d+n)=\alpha(2^{L}-2)=L-1=(r+n-3)-1=r+n-4.$
Since $n=2^{r}+1$, we have
$\alpha(d+n)=r+2^{r}-3.$
We compare this with $n=2^{r}+1$:
$\alpha(d+n)>n\quad\Longleftrightarrow\quad r+2^{r}-3>2^{r}+1\quad\Longleftrightarrow\quad r-3>1\quad\Longleftrightarrow\quad r>4.$
By hypothesis $r\geq 5$, so indeed $\alpha(d+n)>n$. By Proposition 2.1, this is equivalent to $\beta(d)>n$. ∎
Remark 4.5.
For small values of $r$ one can verify the above calculation by direct computation. For example, when $r=5$ we have $n=33$, $k=29$ and $d+n=2^{35}-2$, whose binary expansion has $34$ ones, so $\alpha(d+n)=34>33=n$. For $r=6$ we obtain $n=65$, $k=61$ and $d+n=2^{68}-2$ with $\alpha(d+n)=67>65$, and so on.
Theorem 4.4 gives an infinite family of pairs $(n,k)$ with k=n4k=n-4 for which $\beta(d)>n$. In the next section we combine this with the odd-parity non-hit theorem to produce new motivic Peterson counterexamples.
5. Base change and motivic counterexamples over general fields

### v4

@@ 3.2. A linear algebra lemma
3.3. The local $Q_{0}$–image in the $M_{1}$–summand
We now translate Kameko’s description of the image of $Q_{0}$ into a parity statement for the local map $\vartheta$.
@@ 3.3. The local $Q_{0}$–image in the $M_{1}$–summand
    The set $\pi(M_{0}\cup M_{1})$ is a basis of
    $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
    In particular,
    $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V.$
@@
    For each monomial $z\in N_{n}^{d-1,*}$ with $\omega_{0}(z)=k+1$, the element
    $\pi\bigl(Q_{0}(z)\bigr)\in\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
    is a linear combination of $\pi(z^{\prime})$ with $z^{\prime}\in M_{0}$ and elements of the form $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$ with $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$.
@@
    There exist monotone injections $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ (explicitly, the ones with $\sigma_{1}(1)=1,\sigma_{1}(2)=3,\dots,\sigma_{1}(k)=k+1$ and $\sigma_{2}(1)=2,\dots,\sigma_{2}(k)=k+1$) and a monomial $z\in N_{n}^{d-1,*}$ with $(\omega_{0}(z),\omega_{1}(z))=(k+1,n-2)$ and $\alpha_{i}(z)=k$ for all $i$, such that
    $p_{M_{1}}\pi\bigl(Q_{0}(z)\bigr)=\pi\bigl(\sigma_{1}(z_{k})+\sigma_{2}(z_{k})\bigr).$
We can now identify the local $Q_{0}$–image in $V$.
Proposition 3.5 (Parity description of the local $Q_{0}$–image).
With notation as above, identify $V$ with $W=\mathbb{F}_{2}^{N}$ by sending the basis $\pi(\sigma_{i}(z_{k}))$ to $e_{i}$. Then
$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)=\ker(\varepsilon)\subset V.$
Equivalently, the local $M_{1}$–part of the $Q_{0}$–image is exactly the even-parity hyperplane.
@@
If $z\in N_{n}^{d-1,*}$ has $\omega_{0}(z)\neq k+1$, then $Q_{0}(z)$ lies in a summand $\Lambda_{n}^{a}\otimes Y_{n}^{2b}$ with $a\neq k$, hence $\operatorname{pr}_{k}(Q_{0}(z))=0$ and therefore $\vartheta(Q_{0}(z))=0$. Thus only monomials with $\omega_{0}(z)=k+1$ contribute.
For such $z$, Proposition 3.4(2) shows that $\pi(Q_{0}(z))$ is a linear combination of $\pi(M_{0})$ and pairwise sums $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$. After applying $p_{M_{1}}$, we conclude that $\vartheta(Q_{0}(z))$ is a linear combination of vectors of the form $e_{i}+e_{j}$. Hence
$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)\subseteq E,$
where $E\subset W$ is the span of all $e_{i}+e_{j}$.
Conversely, Proposition 3.4(3) provides one explicit edge $e_{i_{0}}+e_{j_{0}}$ in the image. Let $S_{n}$ act on $N_{n}$ by permuting the variables $x_{1},\dots,x_{n},y_{1},\dots,y_{n}$. This action commutes with $Q_{0}$, preserves the subspace $G_{n}\subset Y_{n}$, and permutes the basis elements $\pi(\sigma(z_{k}))$ of $V$. Moreover, $S_{n}$ preserves $M_{0}$, since the defining condition for membership in $M_{0}$ is that $\alpha_{i}(z)<k$ for some $i$, and this condition is invariant under permutations of the indices. Hence $U_{0}=\langle\pi(m)\mid m\in M_{0}\rangle$ is $S_{n}$–stable, so the projection
$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
with kernel $U_{0}$ is $S_{n}$–equivariant. Hence, for every permutation $\tau\in S_{n}$,
$\vartheta\bigl(Q_{0}(\tau\cdot z)\bigr)=p_{M_{1}}\pi\operatorname{pr}_{k}\bigl(Q_{0}(\tau\cdot z)\bigr)=\tau\cdot p_{M_{1}}\pi\operatorname{pr}_{k}\bigl(Q_{0}(z)\bigr)=\tau\cdot\vartheta\bigl(Q_{0}(z)\bigr).$
Therefore all edges of the Johnson graph $J(n,k)$ occur in the image, because the orbit of the explicit pair $(\sigma_{1},\sigma_{2})$ under $S_{n}$ consists exactly of pairs of kk–subsets differing in one element.
Since the Johnson graph $J(n,k)$ is connected for $1\leq k\leq n-1$, any vector $e_{i}+e_{j}$ is a sum of edge vectors along a path from $i$ to $j$. Thus every pairwise sum $e_{i}+e_{j}$ lies in $\vartheta(Q_{0}(N_{n}^{d-1,*}))$, so
$E\subseteq\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
Combining both inclusions gives
$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)=E.$
Finally, by Lemma 3.3, $E=\ker(\varepsilon)$. ∎
3.4. From $Q_{0}$ to the full hit image in the local top layer
@@
We now extend Proposition 3.5 from the Bockstein to the full positive-degree part of the motivic Steenrod algebra.
Lemma 3.6 (Reduced-power words vanish under $\vartheta$).
Let $\mathcal{W}$ be the set of finite words in the generators $Q_{0},\mathcal{P}^{1},\mathcal{P}^{2},\dots$ of $A^{\sharp}$. Then $A^{\sharp}_{+}$ is spanned by the images of elements of $\mathcal{W}$. If $w\in\mathcal{W}$ contains some reduced power $\mathcal{P}^{c}$ with $c\geq 1$, then
$\vartheta(w\cdot u)=0$
for every $u\in N_{n}$ such that $w\cdot u\in N_{n}^{d,*}$. Consequently,
$\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
@@
Since $A^{\sharp}_{+}$ is spanned by the images of words in $Q_{0},\mathcal{P}^{1},\mathcal{P}^{2},\dots$, it is enough by linearity to consider an operator represented by a single word $w$.
Assume that $w$ contains at least one reduced power. The reduced powers act trivially on the exterior factor $\Lambda_{n}$ and act on $Y_{n}$ through the reduced-power subalgebra $P\subset A^{\sharp}$. Because the word $w$ contains a reduced power, the induced operator on the $Y_{n}$–factor lies in $P^{+}$. Hence, for every $u\in N_{n}$, the $Y_{n}$–part of $w\cdot u$ belongs to
$P^{+}(Y_{n})\subset G_{n}.$
The Bockstein $Q_{0}$ acts only on the exterior factor and does not alter the $Y_{n}$–part. Therefore, after applying the projection
$\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}},$
the result lies in
$\Lambda_{n}^{k}\otimes G_{n}^{2d_{1}},$
which is annihilated by
$\pi:\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Hence $\vartheta(w\cdot u)=0$.
Finally, since $Q_{0}^{2}=0$, the only nonzero word containing no reduced power is the single letter $Q_{0}$. The final equality follows. ∎
@@
Theorem 3.7 (Theorem 1.4).
Let $n\geq 2$ and $1\leq k<n$, and let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. Let $V$ be the local $M_{1}$–summand defined above, let $\vartheta:N_{n}^{d,*}\to V$ be the local top-layer projection from Definition 3.1, and let $\varepsilon:V\to\mathbb{F}_{2}$ be the parity functional from Definition 3.2. Then:
@@
    $\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
@@
    If $u\in N_{n}^{d,*}$ satisfies $\varepsilon(\vartheta(u))=1$, then $u\notin A^{\sharp}_{+}(N_{n})$.
@@
    In particular, if
    $u=\sum_{\sigma\in S}\sigma(z_{k})\in N_{n}^{d,*}$
    for some non-empty subset $S\subset\mathrm{Mono}(k)$ and $|S|$ is odd, then $u\notin A^{\sharp}_{+}(N_{n})$.
@@
By Lemma 3.6,
$\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
By Proposition 3.5, the latter space is exactly $\ker(\varepsilon)$. This proves (1).
For (2), if $u$ were hit, then $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$, and so part (1) would imply $\vartheta(u)\in\ker(\varepsilon)$, contradicting $\varepsilon(\vartheta(u))=1$.
For (3), note that if $u=\sum_{\sigma\in S}\sigma(z_{k})$, then $u\in\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$, its image under $\pi$ has no $M_{0}$–part, and therefore
$\vartheta(u)=\sum_{\sigma\in S}\pi(\sigma(z_{k})).$
Its parity is exactly $|S|$ modulo $2$. Hence odd cardinality implies $\varepsilon(\vartheta(u))=1$, and the claim follows from (2). ∎
Remark 3.8.
Theorem 3.7 identifies the image of hit elements after applying the local projection $\vartheta$. It does not assert that every even-parity linear combination of the monomials $\sigma(z_{k})$ is itself hit. The statement is weaker and sharper: every even-parity vector occurs as the local $M_{1}$–component of some hit element, while odd-parity vectors cannot occur in that way.
@@
Let $\bar{u}_{S}$ be the image of $u_{S}$ in $QN_{n}^{d,*}(\mathbb{C})$ and hence in $QM_{n}^{d,*}(\mathbb{C})$. If $|S|$ is odd, then $\bar{u}_{S}$ is non-zero. In particular, each individual translate $\sigma(z_{k})$ represents a non-zero class, and any sum of an odd number of translates does as well.
@@
This is Theorem 3.7(3), interpreted in the hit quotient. ∎
4. Base change and motivic Peterson counterexamples over general fields